OpenStudy (anonymous):

What is the sum of the arithmetic sequence 4, 11, 18 …, if there are 26 terms?

hartnn (hartnn):

the first term a1 = 4 common difference d= 11-4 = 18-11 = 7 the sum to n terms (here n=26) is \(\huge S_n= (n/2)(2a+(n-1)d)\) just put the values.

OpenStudy (anonymous):

ok I did that and got 2379. is that right?

hartnn (hartnn):

yup, thats correct!

OpenStudy (anonymous):

ty :))

hartnn (hartnn):

welcome :)